Circular No. 7189

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7244/7440/7444 (for emergency use only)

BL LACERTAE J. R. Mattox, Boston University, reports on behalf of the Whole Earth

Blazar Telescope (WEBT) Collaboration: “BL Lac is currently observed to be bright (G. Tosti measured R = 13.10 on May 28, and T. Balonek measured R = 13.03 on June 4.3 UT). The WEBT will attempt to obtain 24-hr/day optical coverage of BL Lac during two intervals: June 5.330– 7.613, in conjunction with a continuous x-ray observation by BeppoSax; and June 28.226–30.694, in conjunction with a continuous x-ray observation by ASCA. Updated information will be maintained at http://gamma.bu.edu/ webt/bllac 99.html during these campaigns. The participation of additional optical observers is required to obtain continuous coverage.”

SUPERNOVAE 1987T, 1988ah, 1990al J. Mueller reports her discovery of three apparent supernovae on plates

taken with the 1.2-m Oschin Schmidt telescope in the course of the second Palomar Sky Survey:

SN  UT  α_{2000}  δ_{2000}  Mag. Offset
1987T 1987 Sept. 28 21^{h}45^{m}33.^{s}33 +32^{o}06′22′′.8 17.5  9′′ E, 1′′ S
1988ah 1988 Dec. 11 8 08 28.30 +11 10 34.7 16.5  10′′ W, 6′′ S
1990al 1990 July 19 19 09 56.64 +42 31 41.7 16.5  12′′ E, 15′′ N

SN 1987T was confirmed at mag ∼ 17 on a sky-survey plate taken on 1987 Aug. 27 (both IIIa-F plates of SN 1987T were taken by Mueller); no object is present at this position on the Digitized Sky Survey (DSS) or on a skysurvey plate taken on 1989 July 8 by Mueller and J. D. Mendenhall. SN 1988ah in UGC 4244 was confirmed (mag ∼ 18) by B. Skiff, C. S. Shoemaker, and Mueller on hypered 4415 films taken on 1989 Jan. 13 with the 0.46-m Palomar Schmidt telescope by E. M. and C. S. Shoemaker; the IIIa-F discovery plate was taken by C. Brewer and Mueller. No object appears at the location of SN 1988ah on the DSS or on a sky-survey plate taken on 1992 Nov. 22 by Mendenhall and Mueller. SN 1990al was discovered on a IIIa-J plate taken by Brewer, and it also appears (mag ∼ 18) on a IIIa-F plate taken on 1990 Aug. 23 by Mueller and Brewer; nothing appears at this position on the DSS or on a sky-survey plate taken on 1989 July 4 by Brewer and Mendenhall.

1999 June 4  \begin{matrix}c \\ ◯ \end{matrix} Copyright 1999 CBAT  Daniel W. E. Green